OpenStudy (anonymous):

HELP PLEASE : Solve 4x - 2x - 2 = -42. HELP PLEASE : Solve 4x - 2x - 2 = -42. @Mathematics

OpenStudy (anonymous):

ANSWER KEY: -20 -7 -5 0

OpenStudy (zarkon):

-20

OpenStudy (anonymous):

You need to get all terms with x together and all constants together. Then solve for x. 4x - 2x - 2 = -42 => 4x - 2x = -42 + 2 => 2x = -40 => x = -20

OpenStudy (anonymous):

thank you bolth :))